two angles are supplements. The larger angle is 30 degrees more than 3 times the smaller. Find the measures of the two angles.
Answer by checkley79(3341)

X+Y=180
X=3Y+30
(3Y+30)+Y=180
4Y=180-30
4Y=150
Y=150/4
Y=37.5 FOR THE SMALLER ANGLE.
X=3*37.5+30
X=112.5+30
X=142.5 FOR THE LARGER ANGLE.
PROOF:
37.5+142.5=180
180=180